What is the present value of $ 7 comma 000 received a. 14 years from today when the interest rate is 4 % per​ year? b. 28 years from today when the interest rate is 8 % per​ year? c. 7 years from today when the interest rate is 2 % per​ year?
Answer #1

a. Present value = $7,000 / 1.0414 = $4,042.33

b. Present value = $7,000 / 1.0828 = $811.40

c. Present value = $7,000 / 1.027 = $6,093.92